OpenStudy (anonymous):

1. Which two points did you use to draw the line of best fit? Point #1: 59 Point #2: 71 2. Write the equation of the line passing through those two points using the point-slope formula y - y1 = m(x - x1). Show all of your work. Remember to find the slope of the line first. 3. What does the slope of the line represent within the context of your graph? 4. Using the equation that you found in question 2, approximately how tall is a person whose arm span is 66 inches? 5. According to your line of best fit, what is the arm span of a 74-inch-tall person? @agent0smith

OpenStudy (anonymous):

I have no idea how to write an equation for a point slope formula..

OpenStudy (agent0smith):

Post a screenshot

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i sucked at this...

OpenStudy (agent0smith):

was there a table of points, too? screenshot that

OpenStudy (anonymous):

umm.... i dont think there was

OpenStudy (agent0smith):

How'd you make the graph...?

OpenStudy (anonymous):

oh - well all the numbers on it already are ones from the points table...

OpenStudy (agent0smith):

that's what you should screenshot

OpenStudy (anonymous):

i dont have them anymore tho xD i got rid of them after i made the graph lol

OpenStudy (agent0smith):

lol okay then just pick two points on the graph, which is gonna have to be an estimate since you don't have the points

OpenStudy (agent0smith):

(61, 58) and (68, 65)

OpenStudy (anonymous):

oh ok... 61 and 58

OpenStudy (anonymous):

but idk how to make that an equation

OpenStudy (agent0smith):

No those are the two points you picked. Points are coordinates, (x, y)

OpenStudy (agent0smith):

(61, 58) and (68, 65) so from these two find the slope \[\large m = \frac{ y_2 - y_1 }{ x_2 - x_1 }\]

OpenStudy (anonymous):

then what happens?

OpenStudy (agent0smith):

plug in the numbers \[\large m = \frac{ 65-58 }{ 68-61}\]

OpenStudy (anonymous):

ooooooooooo

OpenStudy (anonymous):

i got it - then what? haha

OpenStudy (agent0smith):

What did you get for m?

OpenStudy (anonymous):

m = 13/7

OpenStudy (agent0smith):

Try again :P

OpenStudy (agent0smith):

How'd you get 13?

OpenStudy (anonymous):

ooo xD hahaha omg idk how that happened 7/7

OpenStudy (anonymous):

and that equals 1...

OpenStudy (agent0smith):

lol good. Now use \[\large y - y_1 = m(x - x_1)\]and plug in the slope you found, and one of the points, either (61, 58) or (68, 65)